OpenStudy (anonymous):

What is the smallest value of \[x^2 + 4xy + 5y^2 - 4x - 6y + 7\]

OpenStudy (jamesj):

Use the same method. Complete the squares.

OpenStudy (jamesj):

But this requires a little more thoughtfulness as to what squares can (or should) be completed.

OpenStudy (anonymous):

it does :( thinking time :-D

OpenStudy (anonymous):

I'm sure there's an easy way out of this one if I think hard enough. This is from Spivak

OpenStudy (anonymous):

and it's still the first chapter

OpenStudy (anonymous):

if I can do this one I will be a master at completing squares :-D.... but this might take half an hour :(

OpenStudy (anonymous):

Spivak is really expecting the reader to make a huge leap of intuition here :(

myininaya (myininaya):

i don't like that xy term in there :( we could have just used our algebra skills

myininaya (myininaya):

you know write it in that form where you find the center of the ellipse and then look at the up and down radius to figure out the lowest point

OpenStudy (anonymous):

but this is just in the beginning of the book, where I'm not supposed to know what an ellipse is yet.

myininaya (myininaya):

oh beginning of what book?

OpenStudy (anonymous):

Calculus 3rd, Michael Spivak

myininaya (myininaya):

i thought they talked about ellipse in algebra so this book assumes you don't know the standard form for an ellipse

OpenStudy (anonymous):

I don't really know what it's assuming out of me :-P

myininaya (myininaya):

lol there is no way to tell i don't think

OpenStudy (anonymous):

but this is a chapter about the basic laws of addition and multiplication, so we have to apply those somewhere here.

OpenStudy (jamesj):

You can use whatever you like about geometry. This book is not assuming you've never seen mathematics before.

myininaya (myininaya):

lol

OpenStudy (anonymous):

I think the only geometry in the chapter was the part about absolute values.... so how do we do the ellipse trick here?

OpenStudy (anonymous):

we can't use differentiation here since that's chapter 6 or something

myininaya (myininaya):

like i said i can't remember with that xy term i will be right back i have to get the clothes out of the dryer

myininaya (myininaya):

brb

OpenStudy (anonymous):

\[x^2 + 4xy + 5y^2 - 4x - 6y + 7\]\[=x^2 + 4x(y-1) + 5y^2 - 6y + 7\]\[=x^2 + 4x(y-1) + 4(y-1)^2 + 5y^2 - 6y + 7 - 4y^2 +8y - 4\]\[=(x+2(y-1))^2 + y^2 + 2y + 3\]\[=(x+2(y-1))^2+y^2 + 2y + 1 + 3 -1\]\[=(x+2(y-1))^2 + (y+1)^2 +2\] when x is -4 and y is -1 it seems the expression is equal to 2

OpenStudy (jamesj):

when x = 4, y = -1

OpenStudy (anonymous):

oops when x is 4

OpenStudy (jamesj):

you've got it.

OpenStudy (jamesj):

now do the problem before this one by yourself

OpenStudy (anonymous):

I did it a little earlier.

OpenStudy (jamesj):

good

OpenStudy (anonymous):

I think the answer was -9/4 something